Calculate the power being consumed by a 100Ω resistive element connected across a 240V

supply.


1
Expert's answer
2021-02-15T17:03:25-0500

Let's first find the current flowing through the resistive element from the Ohm's Law:


"I=\dfrac{V}{R}=\dfrac{240\ V}{100\ \Omega}=2.4\ A."

Finally, we can find the power being consumed by a 100Ω resistive element:


"P=I^2R=(2.4\ A)^2\cdot100\ \Omega=576\ W."